When dropped, a ball rebounds to 4/25 of its original height. How high will the ball rebound after the 6th bounce if it is dropped from the height of 25ft?
Answer by Alan3354(69443)

When dropped, a ball rebounds to 4/25 of its original height. How high will the ball rebound after the 6th bounce if it is dropped from the height of 25ft?
----------
h = 25*(4/25)^6
= 4^6/25^5
= 4096/9765625
= 4.194304E-4 feet
= 0.0004194304 feet
= 0.1278 mm
It's no longer bouncing